OpenStudy (anonymous):

15x^7/9y^5 divided by 45y^2/5x^3

OpenStudy (anonymous):

Can you do another beautiful drawing?

OpenStudy (anonymous):

a) 15x^4/y^3 b) 5x^10/27y^7 c) 27y^7/5x^10 d) y^3/15x^4

OpenStudy (anonymous):

\[\frac{ 15x ^{7} }{ 9y ^{5} }\div \frac{ 45y ^{2} }{ 5x ^{3} } = \frac{ 15x ^{7} }{ 9y ^{5} }*\frac{ 5x ^{3} }{ 45y ^{2} } = \frac{ 75x ^{10} }{ 405y ^{7} }=\frac{ 5x ^{10} }{ 27y ^{7} } \]

OpenStudy (anonymous):

Second choice :)

OpenStudy (anonymous):

It is B.

OpenStudy (anonymous):

I second Falls' work :P

OpenStudy (anonymous):

Lol